# Attacking LLM Watermarks by Exploiting Their Strengths

### v1

@@ Appendix B Proof of Theorem 1
Appendix C Validation of Theorem 1
@@
In this section, we validate Theorem 1 by using watermarked texts of varying lengths $l$ and z-scores $z$ to study the relationship between $\frac{s}{l}$ and $\frac{z^{t}-T^{2}}{T^{2}}$ of Unigram watermark. The results are shown in Fig. 5. As anticipated, 85.78% of the maximum allowable tokens to be inserted into the watermarked content satisfy Eq. 4. Given that this equation analyzes expected $s/l$, a small portion of outliers is reasonable. We primarily visualize this result for Unigram due to its clean robustness guarantee. Other watermarks can also reach similar conclusions, but their bounds on $s$ are either complex [KGW+23a] or lack a closed form [KTHL23], making them difficult to visualize. Our empirical findings in Fig. 1 sufficiently prove an attacker can insert nontrivial portions of toxic or incorrect tokens into the watermarked text to launch the spoofing attack, which can be generalized across all robust watermarking schemes.
@@
Figure 5: The relationship between $s/l$ and $z$. The data points are evaluated on Unigram using LLAMA-2-7B and 500500500 samples from OpenGen dataset.
Appendix D Proof of Theorem 2
@@
Recall that KGW and Unigram randomly split the tokens in the vocabulary into the green list and the red list. We consider greedy sampling, where the token with the highest (watermarked) probability is sampled. We have $n$ independent observations under different watermark keys. For each key, the token $\textbf{x}_{i}$ with the highest unwatermarked probability in the green list is $\gamma$. As long as $\textbf{x}_{i}$ is the green list, the greedy sampling will always yield $\textbf{x}_{i}$ since the watermarks add the same constant to all the tokens’ logits in the green list.
@@
Consider we have $c$ such tokens having the potential to exceed $\textbf{x}_{i}$. Then at least one of the $c$ tokens is in the green list for at least $k$ keys among the other $n-k$ keys:
@@
Thus, with all the above analysis, we have that if there are $c$ tokens that have the potential to exceed the probability of the token with the highest unwatermarked probability (i.e., $\textbf{x}_{i}$), the probability that the most frequent token among the $n$ observations is the same as $\textbf{x}_{i}$ is:
@@
Appendix E Probability Bound of Unwatermarked Token Estimation for Exp
In this section, we present and prove the probability bound of unwatermarked token estimation for the Exp watermark [KTHL23].
@@ Appendix E Probability Bound of Unwatermarked Token Estimation for Exp
Suppose there are $n$ observations under different keys, the highest probability for the unwatermarked tokens is $p$. Then the probability that the most frequently appeared token among the $n$ observations is the same as the original unwatermarked token with the highest probability is:
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}p^{k}(1-p)^{n-k}$
@@
The proof of Theorem 3 is straightforward. As we have introduced in Appendix A, the Exp watermark employs the Gumbel trick sampling [Gum48] when embedding the watermark. Thus, the probability that we observe the token whose original unwatermarked probability is $p$ is exactly $p$ for each of the independent keys. Thus, if we make $n$ observations under different keys, then at least half of them yield the token with the highest original probability $p$:
@@
Appendix F Algorithms of Attacks Exploiting the Detection API
In this section, we provide the detailed algorithm of the attacks exploiting the detection API as we have introduced in Sec. 6. Specifically, we present the algorithm for watermark-removal attack exploiting the detection API in Alg. 1 and the algorithm for spoofing attack exploiting the detection API in Alg. 2.
Algorithm 1 Watermark-removal attack exploiting the detection API.
  Input: Prompt $\textbf{x}_{\text{prompt}}$, watermarked LLM $M_{\text{wm}}$, detection API $f_{\text{detection}}$, maximum output token number $m\geq 2$
  Let $k\leftarrow 5$, $\textbf{x}_{1}\sim M_{\text{wm}}(\textbf{x}_{\text{prompt}})$
  for $t=2$ to $m$ do
  $(\textbf{x}_{t}^{1},\textbf{x}_{t}^{2},\cdots,\textbf{x}_{t}^{k}),(\textbf{p}_{t}^{1},\textbf{p}_{t}^{2},\cdots,\textbf{p}_{t}^{k})\leftarrow M_{\text{wm}}(\textbf{x}_{prompt}||\textbf{x}_{1}\cdots\textbf{x}_{t-1})$The watermarked LLM returns the top $k$ tokens and their corresponding probabilities in descending order.}
  for $i=1$ to $k$ do
   $d_{i}\leftarrow f_{\text{detection}}(\textbf{x}_{1}||\cdots||\textbf{x}_{t-1}||\textbf{x}_{t}^{i})$
  $d_{\text{min}}\leftarrow\min(d_{1},d_{2},\cdots,d_{k})$, $l_{\text{candidate}}\leftarrow\text{empty}$Get the detection score with the lowest confidence.}
  for $i=1$ to $k$ do
   if $d_{\text{min}}=d_{i}$ then
    $l_{\text{candidate}}\leftarrow l_{\text{candidate}}||\textbf{x}_{t}^{i}$Get all the tokens with the lowest detection confidence.}
  if $\textbf{x}_{t}^{1}\in l_{\text{candidate}}$ then
   $j\leftarrow 0$If the token with the highest probability (the first token) is in the list, output that token.}
  else
   $c\leftarrow 1$
   for $\textbf{x}_{t}^{i}\in l_{\text{candidate}}$ do
    $\textbf{p}_{t}^{i}\leftarrow\textbf{p}_{t}^{1}/c$Update the tokens’ probabilities that have lowest detection confidence scores.}
    $c\leftarrow c+1$
   $\textbf{p}_{t}^{1}\leftarrow 0$
   $j\leftarrow\text{Sample}(\textbf{p}_{t}^{1},\cdots,\textbf{p}_{t}^{k})$Sample the tokens according to the updated probabilities.}
  $\textbf{x}_{t}\leftarrow\textbf{x}_{t}^{j}$
  Return $\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{m}$
Algorithm 2 Spoofing attack exploiting the detection API.
  Input: Prompt $\textbf{x}_{\text{prompt}}$, local LLM $M$, detection API $f_{\text{detection}}$, maximum output token number $m$
  Let $k\leftarrow 3$
  for $t=1$ to $m$ do
  $(\textbf{x}_{t}^{1},\textbf{x}_{t}^{2},\cdots,\textbf{x}_{t}^{k}),(\textbf{p}_{t}^{1},\textbf{p}_{t}^{2},\cdots,\textbf{p}_{t}^{k})\leftarrow M(\textbf{x}_{prompt}||\textbf{x}_{1}\cdots\textbf{x}_{t-1})$The local LLM returns the top $k$ tokens and their corresponding probabilities in descending order.}
  for $i=1$ to $k$ do
   $d_{i}\leftarrow f_{\text{detection}}(\textbf{x}_{1}||\cdots||\textbf{x}_{t-1}||\textbf{x}_{t}^{i})$
  $j\leftarrow\arg\max(d_{1},d_{2},\cdots,d_{k})$Get the token resulting in the highest confidence.}
  $\textbf{x}_{t}\leftarrow\textbf{x}_{t}^{j}$
  Return $\textbf{x}_{1},\textbf{x}_{2},\cdots,\textbf{x}_{m}$
@@ Appendix F Algorithms of Attacks Exploiting the Detection API
Appendix G Additional Results of Watermark-Removal Attacks Exploiting Quality-Preserving Property
@@
In this section, we provide more evaluation results of the watermark-removal attack exploiting the quality-preserving property (see Sec. 5) on all the three watermarks (KGW, Unigram, and Exp) and two models (Llama-2-7B and OPT-1.3B). The results are shown in Fig. 678910. For KGW watermark on OPT-1.3B model and Unigram watermark on Llama-2-7B and OPT-1.3B models, we have consistent observations with the KGW watermark on Llama-2-7B as we present in Sec. 5.1, demonstrating the effectiveness of our attacks. For the Exp watermark, our results in Fig. 7 and Fig. 10 also show the watermark can be easily removed using multiple queries to estimate the unwatermarked tokens.
@@
For Exp watermark [KTHL23], the use of multiple watermark keys is inherent in its design, which is default to 256256256. Thus, Exp has the same unwatermarked detection results for various numbers of queries under different watermark keys. From the results in Fig. 7 and Fig. 10, we conclude that using $n=13$ queries under different keys, the resulting p-value is very close to that of the content without watermark and is significantly different from the watermarked p-value, which shows that we can effectively remove the watermark using 131313 queries for each token. We note that for Exp, the perplexity of watermarked content is significantly higher than that of unwatermarked content. This is primarily because Exp does not allow sampling in watermark embedding, which becomes a deterministic algorithm when the key is fixed. Conversely, our watermark-removal attack generates content with much lower perplexity, making it comparable to unwatermarked content when the query number under different keys exceeds 131313. This can be attributed to our attack functioning as a layer of random sampling. Unlike greedy sampling methods, we have a probability to sample the token with the highest unwatermarked probability (refer Sec. 4, Appendix D, and Appendix E).
The results of the three watermarks and two models prove that the watermark-removal attack exploiting the quality-preserving property using multiple keys can effectively eliminate the watermarks while maintaining high output quality. We anticipate that this attack can be generalized to all watermarking schemes that have quality-preserving properties.
(a) Z-Score.
(b) Perplexity.
@@
Figure 6: Watermark-removal on Unigram watermark [ZALW24] and Llama-2-7B model with multiple watermark keys.
(a) P-Value.
(b) Perplexity.
@@
Figure 7: Watermark-removal on Exp watermark [KTHL23] and Llama-2-7B model with multiple watermark keys.
(a) Z-Score.
(b) Perplexity.
@@
Figure 8: Watermark-removal on KGW watermark [KGW+23a] and OPT-1.3B model with multiple watermark keys.
(a) Z-Score.
(b) Perplexity.
@@
Figure 9: Watermark-removal on Unigram watermark [ZALW24] and OPT-1.3B model with multiple watermark keys.
(a) P-Value.
(b) Perplexity.
Figure 10: Watermark-removal on Exp watermark [KTHL23] and OPT-1.3B model with multiple watermark keys.
Appendix H Additional Results of Attacks Exploiting Detection APIs
[…]
(a) Z-Score/P-Value of removal.
(b) Perplexity of removal.
@@ Appendix H Additional Results of Attacks Exploiting Detection APIs
Appendix I Additional Results of DP Defense
We present additional evaluation results of our defense technique that enhances the watermark detection by utilizing the techniques of differential privacy (see Sec. 6). Consistent with Sec. 6.3, we evaluate the utility of the DP defense as well as its performance in mitigating the spoofing attack exploiting the detection API. The results are shown in Fig. 1213141516.
[…]
Our defense can be generalized to all LLM watermarking schemes. It allows us to substantially mitigate spoofing attacks exploiting the detection API while having a negligible impact on utility.
@@ Appendix I Additional Results of DP Defense
Figure 12: Evaluation of DP watermark detection on Unigram watermark and Llama-2-7B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).
@@
Figure 13: Evaluation of DP watermark detection on Exp watermark and Llama-2-7B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP under different query numbers. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) Z-scores with/without DP and under multiple queries.

### v3

@@ Appendix D Attack Feasibility Analysis of Piggyback Spoofing Exploiting Robustness
Different from the analysis in the Unigram paper on how the z-score changes given a specific number of edits, we have a tight bound on the maximum possible number of edits, which is also more straightforward for the attack feasibility analysis. According to Theorem 1, as long as the number of toxic tokens inserted is bounded by $l\frac{z^{2}-T^{2}}{T^{2}}$, the attacker can execute a piggyback attack to generate toxic content with the target watermark embedded. The editing distance bound (Def. 3) for a sentence is $\epsilon=l\frac{z^{2}-T^{2}}{T^{2}}$. A stronger watermark makes piggyback spoofing attacks easier by allowing more toxic tokens to be inserted. This conclusion applies universally to all robust watermarking schemes. This is a fundamental design trade-off: if a watermark is robust, such spoofing attacks are inevitable and may be extremely difficult to detect, as even one toxic token can render the entire content harmful or inaccurate.
Appendix E Validation of Theorem 1
@@
In this section, we validate Theorem 1 by using watermarked texts of varying lengths $l$ and z-scores $z$ to study the relationship between $\frac{s}{l}$ and $\frac{z^{t}-T^{2}}{T^{2}}$ of Unigram watermark. The results are shown in Fig. 3(e). As anticipated, 85.78% of the maximum allowable tokens to be inserted into the watermarked content satisfy Theorem 1. Given that this equation analyzes expected $s/l$, a small portion of outliers is reasonable. We primarily visualize this result for Unigram due to its clean robustness guarantee. Other watermarks can also reach similar conclusions, but their bounds on $s$ are either complex [17] or lack a closed form [20], making them difficult to visualize. Our empirical findings in Fig. 1 sufficiently prove an attacker can insert nontrivial portions of toxic or incorrect tokens into the watermarked text to launch the spoofing attack, which can be generalized across all robust watermarking schemes.
@@
Figure 5: The relationship between $s/l$ and $z$. The data points are evaluated on Unigram using LLAMA-2-7B and $500$ samples from OpenGen dataset.
Appendix F Piggyback Attack Examples
Here we present more piggyback attack results using the edition strategy by querying GPT4 using the prompt “Modify less than 3 words in the following sentence and make it inaccurate or have opposite meanings.” The attack is launched on KGW watermark and LLAMA-2-7B model.
Watermarked content, z-score: $4.93$, PPL: $4.61$ Earth has a history of 4.5 billion years and humans have been around for 200,000 years. Yet humans have been using computers for just over 70 years and even then the term was first used in 1945. In the age of technology, we are still just getting started. The first computer, ENIAC (Electronic Numerical Integrator And Calculator), was built at the University of Pennsylvania between 1943 and 1946. The ENIAC took up 1800 sq ft and had 18,000 vacuum tube and mechanical parts. The ENIAC was used for mathematical calculations, ballistics, and code breaking. The ENIAC was 1000 times faster than any other calculator of the time. The first computer to run a program was the Z3, built by Konrad Zuse at his house.
Piggyback attack, z-score: $4.36$, PPL: $5.68$ Earth has a history of 4.5 billion years and humans have been around for 200,000 years. Yet humans have been using computers for just over 700 years and even then the term was first used in 1445. In the age of technology, we are still just getting started. The first computer, ENIAC (Electronic Numerical Integrator And Calculator), was built at the University of Pennsylvania between 1943 and 1946. The ENIAC took up 1800 sq ft and had 18,000 vacuum tube and mechanical parts. The ENIAC was used for mathematical calculations, ballistics, and code breaking. The ENIAC was 1000 times slower than any other calculator of the time. The first computer to run a program was the Z3, built by Konrad Zuse at his house.
Watermarked content, z-score: $4.57$, PPL: $3.51$ Harry Potter is a 2001 film adaptation of J.K. Rowling’s best-selling novels. In the year 1991, Harry Potter is rescued from his aunt and uncle’s home and taken to Hogwarts, a school for magic. During the course of his first year, Harry makes friends and learns about the evil Lord Voldemort and his attempts to kill him. The film was directed by Chris Columbus, and stars Daniel Radcliffe as Harry Potter, with Richard Harris, Maggie Smith, Robbie Coltrane, Alan Rickman, Emma Watson, Rupert Grint, Gary Oldman, and Julie Walters in supporting roles. In the first film, Harry Potter is taken to Hogwarts School of Witchcraft and Wizardry, where he makes friends and discovers that Lord Voldemort wants to kill him.
Piggyback attack, z-score: $4.47$, PPL: $3.78$ Harry Potter is a 2001 film adaptation of J.K. Rowling’s best-selling novels. In the year 2991, Harry Potter is rescued from his aunt and uncle’s home and taken to Hogwarts, a school for magic. During the course of his first year, Harry makes friends and learns about the evil Lord Voldemort and his attempts to kill him. The film was directed by Chris Columbus, and stars Daniel Radcliffe as Harry Potter, with Richard Harris, Maggie Smith, Robbie Coltrane, Alan Rickman, Emma Watson, Rupert Grint, Gary Oldman, and Julie Walters in supporting roles. In the first film, Harry Potter is taken to Hogwarts School of Witchcraft and Wizardry, where he makes friends and discovers that Lord Voldemort wants to kill him.
Watermarked content, z-score: $6.15$, PPL: $4.29$ Usain Bolt is a 6’6”, 200-pound sprinter from the island of Jamaica. He’s also arguably the greatest athlete to ever walk the earth. In an era where many of us are glued to our phones, he has over 18 million followers on Instagram. The 31-year-old athlete is known for his incredible speed and endurance. He’s won eight Olympic gold medals, including three in 2012 alone. He’s also set world records in both the 100-meter dash and 200-meter dash. Bolt is the only person to ever hold the world record in both events at the same time, which he did twice. He’s also the only person to ever hold the world record in both events at the same Olympics. Bolt is known for his incredible speed and endurance.
Piggyback attack, z-score: $5.58$, PPL: $4.46$ Usain Bolt is a 6’6”, 200-pound sprinter from the island of Jamaica. He’s also arguably the greatest athlete to ever walk the earth. In an era where many of us are glued to our phones, he has over 18 million followers on Instagram. The 31-year-old athlete is known for his incredible speed and endurance. He’s won three Olympic gold medals, including three in 2012 alone. He’s also set world records in both the 100-meter dash and 200-meter dash. Bolt is the only person to ever hold the world record in both events at the same time, which he did twice. He’s also the only person to ever hold the world record in both events at the same Olympics. Bolt is known for his incredible speed and endurance.
Watermarked content, z-score: $6.01$, PPL: $6.68$ The history of the modern airplane is 100 years old this month. And yet it’s not been 100 years since the Wright Brothers’ first flight. The first airplane flight took place on Dec. 17, 1903. After three years of development, Orville and Wilbur Wright’s first flight lasted only 12 seconds. But within a decade, the first airliner flew. In 1924, the Ford Motor Co. flew the first commercial plane on the U.S. East Coast. In the next year, the company built a 10-passenger airliner with passenger windows and seats and an aisle. The 10-seat plane was called the Model T, and Ford executives said it would have been better if the company made a 10-passenger car instead of a plane.
Piggyback attack, z-score: $5.03$, PPL: $7.19$ The history of the modern airplane is 100 years old this month. And yet it’s not been 100 years since the Wright Brothers’ first flight. The first airplane flight took place on Dec. 17, 1903. After three years of development, Orville and Wilbur Wright’s first flight lasted only 12 seconds. But within a decade, the first airliner flew. In 1924, the Ford Motor Co. never flew the first commercial plane on the U.S. East Coast. In the next year, the company built a 10-passenger airliner with passenger windows and seats and an aisle. The 10-seat plane was called the Model T, and Ford executives said it would have been better if the company made a 10-passenger car instead of a plane.
Appendix G Additional Results of Piggyback Spoofing Attack
In Sec. 4, we present the piggyback spoofing attack using toxic token insertion strategy on LLAMA-2-7B model. Here, we present the results on OPT-1.3B model, which are consistent with LLAMA-2-7B model’s results.
Figure 6: Piggyback spoofing of robust watermarks with toxic token insertion strategy on OPT-1.3B.
In Sec. 4, we present the fluent inaccurate editing strategy by querying the GPT4 on KGW watermark and LLAMA-2-7B model. Here we present more results of this strategy on all the three watermarks (KGW, Unigram, and Exp) and two models (LLAMA-2-7B and OPT-1.3B). The results are shown in Fig. 7, Fig. 8, and Fig. 9, which are consistent with our findings in Fig. 1, indicating that our piggyback spoofing attack can be generalized across various robust watermarks and models.
(a) LLAMA-2-7B model.
(b) OPT-1.3B model.
Figure 7: Fluent inaccurate editing strategy on KGW watermark and LLAMA-2-7B and OPT-1.3B models.
(a) LLAMA-2-7B model.
(b) OPT-1.3B model.
Figure 8: Fluent inaccurate editing strategy on Unigram watermark and LLAMA-2-7B and OPT-1.3B models.
(a) LLAMA-2-7B model.
(b) OPT-1.3B model.
Figure 9: Fluent inaccurate editing strategy on Exp watermark and LLAMA-2-7B and OPT-1.3B models.
Appendix H Watermark Key Number Analysis for Watermark-Removal Attacks Exploiting the Use of Multiple Watermark Keys
Now we analyze the number of required queries under different keys to estimate the token with the highest probability without a watermark. We have the following probability bound for KGW and Unigram with the corresponding proof, and present the bound for Exp in Appendix I.
Theorem 2 (Probability bound of unwatermarked token estimation).
$\footnotesize 1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}\gamma^{k}(1-\gamma)^{n-k}\times p(k),$ (6)
In a practical scenario where $n=13,\gamma=0.5$, and $c=3$, Theorem 2 suggests that the attacker has a probability of $0.71$ in finding the token with the highest unwatermarked probability. This implies that we can successfully remove watermarks from over $71\%$ of tokens using a small number of observations under different keys ($n=13$), yielding high-quality unwatermarked content.
@@ Appendix G Additional Results of Piggyback Spoofing Attack
Recall that KGW and Unigram randomly split the tokens in the vocabulary into the green list and the red list. We consider the greedy sampling, where the token with the highest (watermarked) probability is sampled. We have $n$ independent observations under different watermark keys. For each key, the token $\textbf{x}_{i}$ with the highest unwatermarked probability is in the green list is $\gamma$. As long as $\textbf{x}_{i}$ is the green list, the greedy sampling will always yield $\textbf{x}_{i}$ since the watermarks add the same constant to all the tokens’ loogits in the green list.
@@
Consider we have $c$ such tokens having potential to exceed $\textbf{x}_{i}$. Then at least one of the $c$ tokens is in the green list for at least $k$ keys among the other $n-k$ keys is:
@@
Thus, with all the above analysis, we have that if there are $c$ tokens that have the potential to exceed the probability of the token with highest unwatermarked probability (i.e., $\textbf{x}_{i}$), the probability that the most frequent token among the $n$ observations is the same as $\textbf{x}_{i}$ is:
@@
Here we consider that the watermarked LLM is utilizing greedy sampling. In practice, the greedy sampling might not be an optimal sampling strategy, but we note that it is extremely challenging to incorporate the multinomial sampling when analyzing the KGW and Unigram watermarks. Because KGW and Unigram add bias to the output logits, which will go through the softmax function to calculate the probabilities for the tokens. Given the softmax function is not unbiased, we cannot get a tight bound on its variance. Thus, we leave this part as a future direction to further incorporate multinomial sampling in the analysis. Nevertheless, our empirical results still show that the attackers can generate high-quality unwatermarked content when multinomial sampling is used. Also, our analysis on Exp watermark in Appendix I can naturally incorporate multinomial sampling.
Appendix I Probability Bound of Unwatermarked Token Estimation for Exp
In this section, we present and prove the probability bound of unwatermarked token estimation for the Exp watermark [20].
@@ Appendix I Probability Bound of Unwatermarked Token Estimation for Exp
Suppose there are $n$ observations under different keys, the highest probability for the unwatermarked tokens is $p$. Then the probability that the most frequently appeared token among the $n$ observations is the same as the original unwatermarked token with highest probability is:
$1-\sum_{k=0}^{\lfloor n/2\rfloor}\binom{n}{k}p^{k}(1-p)^{n-k}$ (7)
@@
The proof of Theorem 3 is straightforward. As we have introduced in Appendix C, the Exp watermark employs the Gumbel trick sampling [12] when embedding the watermark. Thus, the probability that we observe the token whose original unwatermarked probability is $p$ is exactly $p$ for each of the independent keys. Thus, if we make $n$ observations under different keys, then at least half of them yields the token with the highest original probability $p$ is:
@@
Appendix J Additional Results of Watermark-Removal Attacks Exploiting the use of Multiple Watermark Keys
@@
In this section, we provide more evaluation results of the watermark stealing [15] and our watermark-removal attacks exploiting the use of multiple watermark keys (see Sec. 5) on all the three watermarks (KGW, Unigram, and Exp) and two models (LLAMA-2-7B and OPT-1.3B). The results are shown in Fig. 11, Fig. 12, Fig. 13, Fig. 14, Fig. 15. For KGW watermark on OPT-1.3B model and Unigram watermark on LLAMA-2-7B and OPT-1.3B models, we have consistent observations with the KGW watermark on LLAMA-2-7B as we present in Sec. 5.1, demonstrating the effectiveness and generalizability of our attacks. For the Exp watermark, our results in Fig. 12 and Fig. 15 also show that the watermark can be easily removed using multiple queries to estimate the distribution of the unwatermarked tokens.
The results of the watermark stealing [15] on Unigram watermark and OPT-1.3B model are also consistent with our observations in Sec. 5. Using more keys can effectively mitigate the watermark stealing; however, it will make the system more vulnerable to our watermark removal attacks. Throughout these experiments, we observe that using three keys is the optimal choice to defend against both attacks. However, the attack success rates with three keys are not negligible. Thus, consistent with our guidelines in Sec. 5, we highly recommend that the LLM service provider to simultaneously limit the ability of the potentially malicious users.
To further verify that the LLM service provider can mitigate the watermark stealing attacks by limiting the attacker’s query rates, we present the stealing attack results with various numbers of queries on the KGW watermark and LLAMA-2-7B model using three keys in Fig. 10. The results show that by limiting the query rates of the attacker, the attack success rate of the watermark stealing attack can be significantly decreased. Thus, we recommend that the LLM service provider follow a “defense-in-depth” approach and utilize complementary techniques such as anomaly detection, query rate limiting, and user identification verification to mitigate stealing and removal attacks.
Figure 10: Watermark stealing attack [15] on KGW watermark and LLAMA-2-7B model using three keys with different numbers of attacker obtained tokens Q (in million). The attack success rates are based on the threshold with FPR@1e-3.
@@
We note that the watermark stealing attacks do not work on the Exp watermark [20], as the use of a large number of watermark keys is inherent in their design, which defaults to $256$. Thus, we omit the watermark stealing results on Exp, but we show that this watermark is inherently vulnerable to our watermark removal attack. From the results in Fig. 12 and Fig. 15, we conclude that using $n=13$ queries, the resulting p-value is very close to that of the content without a watermark and is significantly different from the watermarked p-value, which shows that we can effectively remove the watermark using $13$ queries for each token. We note that for Exp, the perplexity of the watermarked content is significantly higher than that of the unwatermarked content. This is mainly because Exp does not allow sampling in watermark embedding, which becomes a deterministic algorithm when the key is fixed. In contrast, our watermark removal attack generates content with much lower perplexity, making it comparable to unwatermarked content when the query number under different keys exceeds $13$. This can be attributed to our attack functioning as a layer of random sampling. Unlike greedy sampling methods, we have a probability to sample the token with the highest unwatermarked probability (see Sec. 4, Appendix H, and Appendix I). The results of the three watermarks and two models prove that the watermark-removal attack exploiting the use of multiple watermark keys can effectively eliminate the watermarks while maintaining high output quality.
(a) Z-Score and attack success rate (ASR) of watermark stealing [26].
(b) Z-Score and attack success rate (ASR) of watermark-removal.
(c) Perplexity (PPL) of watermark-removal.
@@
Figure 11: Spoofing attack based on watermark stealing [26] and watermark-removal attacks on Unigram watermark and LLAMA-2-7B model with different number of watermark keys $n$.
(a) P-Value of watermark-removal attack.
(b) PPL of watermark-removal attack.
@@
Figure 12: Watermark-removal on Exp watermark [20] and LLAMA-2-7B model with multiple watermark keys.
(a) Z-Score and attack success rate (ASR) of watermark stealing [26].
(b) Z-Score and attack success rate (ASR) of watermark-removal.
(c) Perplexity (PPL) of watermark-removal.
@@
Figure 13: Spoofing attack based on watermark stealing [26] and watermark-removal attacks on KGW watermark and OPT-1.3B model with different number of watermark keys $n$.
(a) Z-Score and attack success rate (ASR) of watermark stealing [26].
(b) Z-Score and attack success rate (ASR) of watermark-removal.
(c) Perplexity (PPL) of watermark-removal.
@@
Figure 14: Spoofing attack based on watermark stealing [26] and watermark-removal attacks on Unigram watermark and OPT-1.3B model with different number of watermark keys $n$.
(a) P-Value of watermark-removal attack.
(b) PPL of watermark-removal attack.
Figure 15: Watermark-removal on Exp watermark [20] and OPT-1.3B model with multiple watermark keys.
Appendix K Additional Results of Attacks Exploiting Detection APIs
We present the results of watermark-removal and spoofing attacks on OPT-1.3B model in Fig. 16 and Table 3. The results are consistent with the LLAMA-2-7B model presented in Sec. 6.1., with all the attack success rates higher than $75\%$ using a small number of queries to the detection API of around $3$ per token. The results on OPT-1.3B model further demonstrate the effectiveness of our attacks exploiting the detection API.
(a) Z-Score/P-Value of wm-removal.
(b) Perplexity of wm-removal.
@@
Appendix L Additional Results of DP Defense
We present additional evaluation results of our defence technique that enhances the watermark detection by utilizing the techniques of differential privacy (see Sec. 6). Consistent with Sec. 6.3, we evaluate the utility of the DP defense as well as its performance in mitigating the spoofing attack exploiting the detection API. The results are shown in Fig. 17, Fig. 18, Fig. 19, Fig. 20, Fig. 21.
We first identify the optimal noise scale parameter $\sigma$ based on its detection accuracy and attack success rate, aiming for a drop in detection accuracy within $2\%$ and the lowest attack success rate. Then we assess the performance of the defense. Our findings across three watermarks and two models consistently demonstrate that we can significantly reduce the attack success rate to around or below $20\%$.
Our defense can be generalized to all LLM watermarking schemes. It allows us to substantially mitigate spoofing attacks exploiting the detection API while having negligible impact on utility.
(a) Detection ACC and spoofing ASR.
(b) Z-scores with/without DP.
Figure 17: Evaluation of DP watermark detection on Unigram watermark and LLAMA-2-7B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) P-values with/without DP and under multiple queries.
Figure 18: Evaluation of DP watermark detection on Exp watermark and LLAMA-2-7B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP. We use the best $\sigma=4$ from (a).
@@
Figure 19: Evaluation of DP watermark detection on KGW watermark and OPT-1.3B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP. We use the best $\sigma=4$ from (a).
(a) Detection accuracy and spoofing attack success rate.
(b) Z-scores with/without DP and under multiple queries.
Figure 20: Evaluation of DP watermark detection on Unigram watermark and OPT-1.3B model. (a). Detection accuracy and spoofing attack success rate without and with DP watermark detection under different noise parameters. (b). Z-scores of original text without attack, spoofing attack without DP, and spoofing attacks with DP. We use the best $\sigma=4$ from (a).